# On De Giorgi’s Conjecture of Nonlocal approximations for free-discontinuity problems: The symmetric gradient case

### v1

@@ 2. Definitions and notation
for every measurable function $u\colon\text{Dom}(u)\to\mathbb{R}$ such that $A\subset\text{Dom}(u)\cap(\text{Dom}(u)-\varepsilon)$, where $\text{Dom}(u)$ denotes the domain of $u$. For the sequel it is convenient to introduce the following class of Mumford-Shah and Griffith functionals. For $\gamma>0$, the former reads as
$\mathcal{MS}_{\gamma}(u,\Omega)=\begin{cases}\gamma\int_{\Omega}|\nabla u(x)|^{2}\textnormal{d}x+\mathcal{H}^{n-1}(J_{u}),\quad u\in\text{GSBV}(\Omega),\\ +\infty,\qquad u\in L^{1}_{loc}(\Omega)\setminus\text{GSBV}(\Omega).\end{cases}$
@@
$\mathcal{G}_{\lambda}(u,\Omega)=\begin{cases}\lambda\int_{\Omega}|e(u)(x)|^{2}\textnormal{d}x+\mathcal{H}^{n-1}(J_{u}),\quad u\in\text{GSBD}(\Omega),\\ \infty,\qquad u\in L^{0}(\Omega;\mathbb{R}^{n})\setminus\text{GSBD}(\Omega).\end{cases}$
@@
Next, we recall a measurability property originally stated in [21, Lemma 3.6].
@@
We say that a $\mathcal{L}^{n}$-measurable function $u\colon\Omega\to\mathbb{R}^{n}$ belongs to the space $\textnormal{GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ if there exists a positive Radon measure $\lambda\in\mathcal{M}^{+}_{b}(\Omega)$ such that for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ and for $\mathcal{H}^{n-1}$-a.e. $y\in\Pi^{\xi}$ the function $\hat{u}_{y}^{\xi}\in\textnormal{BV}_{loc}(\Omega_{y}^{\xi})$ and
$\int_{\mathbb{S}^{n-1}}\bigg{(}\int_{\Pi^{\xi}}|D\hat{u}_{y}^{\xi}|(U_{y}^{\xi}\setminus J^{1}_{\hat{u}_{y}^{\xi}})+\mathcal{H}^{0}(U_{y}^{\xi}\cap J^{1}_{\hat{u}_{y}^{\xi}})\textnormal{d}\mathcal{H}^{n-1}(y)\bigg{)}\textnormal{d}\mathcal{H}^{n-1}(\xi)\leq\lambda(U),$ (2.2)
for every open subset $U$ of $\Omega$. Moreover, we say that $u\in\textnormal{GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ if $u\in\textnormal{GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ and $\hat{u}_{y}^{\xi}\in\textnormal{SBV}_{loc}(\Omega_{y}^{\xi})$ for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ and for $\mathcal{H}^{n-1}$-a.e. $y\in\Pi^{\xi}$.
@@
Let $\Omega\subset\mathbb{R}^{n}$ open and let $u\colon\Omega\to\mathbb{R}^{n}$ be $\mathcal{L}^{n}$-measurable. Then for every $\xi\in\mathbb{S}^{n-1}$ we define the Borel regular measure $\mu^{\xi}_{u}$ in $\Omega$ as follows. We define $\mu^{\xi}_{u}$ on $\Omega$ as the unique Borel regular measure on $\Omega$ that satisfies for every open set $U\subset\Omega$
$\mu^{\xi}_{u}(U):=\int_{\Pi^{\xi}}|D\hat{u}^{\xi}_{y}|(U^{\xi}_{y}\setminus J^{1}_{\hat{u}^{\xi}_{y}})+\mathcal{H}^{0}(U^{\xi}_{y}\cap J^{1}_{\hat{u}^{\xi}_{y}})\,\mathrm{d}\mathcal{H}^{n-1}(y),$
whenever for $\mathcal{H}^{n-1}$-a.e. $y\in\Pi^{\xi}$ the function $\hat{u}^{\xi}_{y}$ belongs to ${\rm BV}_{loc}(U^{\xi}_{y})$, while $\mu^{\xi}_{u}(U)=\infty$ otherwise.
@@
Since the hypothesis of De Giorgi-Letta’s Theorem [4] are fulfilled, the extension of the set function $U\mapsto\mu^{\xi}_{u}(U)$ to a Borel regular measure on $\Omega$ is uniquely guaranteed by the formula
$\mu^{\xi}_{u}(B):=\inf\{\mu^{\xi}_{u}(U):B\subset U,\ U\subset\Omega\text{ open}\}.$
@@
In view of (2.2), if $u\in{\rm GBV}^{\mathcal{E}}(\Omega)$ we have that $\mu^{\xi}_{u}\in\mathcal{M}^{+}_{b}(\Omega)$ for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$. Moreover, for such $\xi$, we can make use of the measurability property contained in Lemma 2.2 to infer
$\mu^{\xi}_{u}(B)=\int_{\Pi^{\xi}}|D\hat{u}^{\xi}_{y}|(B^{\xi}_{y}\setminus J^{1}_{\hat{u}^{\xi}_{y}})+\mathcal{H}^{0}(B^{\xi}_{y}\cap J^{1}_{\hat{u}^{\xi}_{y}})\,\mathrm{d}\mathcal{H}^{n-1}(y),$
@@
For $p\geq 1$, $u\in{\rm GBV}^{\mathcal{E}}(\Omega)$, and $\xi\in\mathbb{S}^{n-1}$ such that $\mu^{\xi}_{u}\in\mathcal{M}^{+}_{b}(\Omega)$, we define $\hat{\mu}^{p}_{u}$ on $\Omega$ as the unique Borel regular measure on $\Omega$ that satisfies
$\hat{\mu}^{p}_{u}(U):=\sup_{\mathscr{B}}\sum_{B\in\mathscr{B}}\bigg{(}\int_{\mathbb{S}^{n-1}}\mu^{\xi}_{u}(B)^{p}\,\mathrm{d}\mathcal{H}^{n-1}(\xi)\bigg{)}^{\frac{1}{p}}\qquad\text{for $U\subset\Omega$ open,}$, (2.3)
@@
Observe that for $p=1$ we have
$\hat{\mu}^{1}_{u}(U)=\int_{\mathbb{S}^{n-1}}\mu^{\xi}_{u}(U)\,\mathrm{d}\mathcal{H}^{n-1}(\xi)\qquad\text{for $U\subset\Omega$ open}.$ (2.4)
@@ 3. Structure of the space $\textnormal{GBV}^{\mathcal{E}}$
The present section addresses the study of the space ${\rm GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ introduced in Definition 2.3. In particular, we focus on the structure of functions $u\in{\rm GSBV}^{\mathcal{E}}(\Omega,\mathbb{R}^{n})$ under the additional assumption $\hat{\mu}^{p}_{u}(\Omega)<\infty$. As a byproduct, we prove the following identification Theorem.
@@
Let $\Omega\subset\mathbb{R}^{n}$ open. A measurable function $u\colon\Omega\to\mathbb{R}^{n}$ belongs to ${\rm GSBD}(\Omega)$ if and only if $u\in{\rm GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ with $\hat{\mu}^{p}_{u}(\Omega)<\infty$ for some $p\in(1,+\infty)$.
@@ 3.1. Preliminaries
We discuss here the measurability of $\xi\mapsto\mu^{\xi}_{u}(B)$ for $B$ Borel subset of $\Omega$.
@@
Whenever $\xi\in\mathbb{S}^{n-1}$ and $U\subset\Omega$ are such that $\mu^{\xi}_{u}(U)<\infty$, equality (3.3) holds in view of [21, Theorem 3.8]. It remains to consider the case $\mu^{\xi}_{u}(U)=\infty$. For simplicity of notation, we denote by $\eta$ the set function given by the right-hand side of (3.3). By contradiction, assume that $\eta(U)<\infty$. In particular, for every $\tau\in\mathcal{T}$ we have
@@
Hence, $D_{\xi}\tau(u\cdot\xi)\in\mathcal{M}_{b}(U)$. By a Carathéodory construction using $\eta$ as Gauge function and the family of all open balls contained in $U$ as set of generators, we obtain a Borel regular measure $\lambda$ on $U$, with $\lambda_{\delta}$ as approximating measure for $\delta>0$ (see, e.g., [24, Section 2.10.1]). For every $V\Subset U$, for every $\delta>0$ small enough, and for every covering $\mathcal{G}$ of $V$ made of open balls in $\mathbb{R}^{n}$ with diameter smaller than $\delta$, we may assume that $\bigcup_{A\in\mathcal{G}}A\subset U$. By Besicovitch covering theorem, there exists a dimensional constant $c(n)>0$ and $\mathcal{G}_{1},\ldots,\mathcal{G}_{c(n)}$ disjoint countable subfamilies of $\mathcal{G}$ such that
@@
In particular, we have that $\lambda_{\delta}(V)\leq c(n)\eta(U)$, which yields the inequality $\lambda(V)\leq c(n)\eta(U)$ for every $V\Subset U$. Taking the limit $V\nearrow U$ we conclude that $\lambda$ is a positive bounded Radon measure on $U$. By using (3.4) it is not difficult to show that actually for every Borel set $B\subset U$ we have that $|D_{\xi}\tau(u\cdot\xi)|(B)\leq\lambda(B)$ whenever $\tau\in\mathcal{T}$. Thus, we are in a position to apply [21, Theorem 3.5] to deduce that $\mu^{\xi}_{u}(U)\leq\lambda(U)<\infty$, which is a contradiction. Thus, it must be $\eta(U)=\infty$ and equality (3.3) is satisfied. ∎
@@
Under the assumptions of Proposition 3.3, for every open set $U\subset\Omega$ the map $\xi\mapsto\mu^{\xi}_{u}(U)$ is lower-semicontinuous on $\mathbb{S}^{n-1}$.
@@
We notice that equality (3.3) holds for every $U\subset\Omega$ open, for every $\xi\in\mathbb{S}^{n-1}$. In particular, the right-hand side of (3.3) is lower-semicontinuous w.r.t. $\xi\in\mathbb{S}^{n-1}$ for fixed $U\subset\Omega$ open, hence $\xi\mapsto\mu^{\xi}_{u}(U)$ is also lower-semicontinuous. ∎
@@
$\hat{\mu}^{1}_{u}\bigg{(}\bigcap_{j=1}^{k}U_{j}\bigg{)}=\int_{\mathbb{S}^{n-1}}\mu^{\xi}_{u}\bigg{(}\bigcap_{j=1}^{k}U_{j}\bigg{)}\,\textnormal{d}\mathcal{H}^{n-1}(\xi)\,.$
Passing to the limit as $k\to\infty$, by dominated convergence (recall that $\mu^{\xi}_{u}(\Omega)\in L^{1}(\mathbb{S}^{n-1})$) we deduce that
$\hat{\mu}^{1}_{u}(K):=\int_{\mathbb{S}^{n-1}}\mu^{\xi}_{u}(K)\,\textnormal{d}\mathcal{H}^{n-1}(\xi)\,.$
@@
In this section we prove that if $u\in\textnormal{GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ then $u\in\textnormal{GSBD}(\Omega)$ under the additional assumption that $\hat{\mu}^{p}_{u}$ is a finite measure for some $p>1$ (the other implication being trivially true). Such implication is a consequence of the following two theorems, concerning the structure properties of functions in ${\rm GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ with $\hat{\mu}^{p}_{u}$ finite.
@@
Let $u\in{\rm GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ and assume that $\hat{\mu}_{u}^{p}$ is a finite measure for some $p>1$. Then for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ we have
@@
$\mu^{\xi}_{u}(B)\leq\lambda(B),\quad\text{for every $B\subset\Omega$ Borel}.$ (3.8)
@@
$\mu^{\xi}_{u}(B)\leq\lambda(B),\quad\text{for every $B\subset\Omega$ Borel}.$ (3.9)
@@
Since $\lambda$ is a finite Borel measure, the above inequality can be further extended to every Borel set $B\subset\Omega$ by outer regularity. This proves (3.8). In order to verify the validity of (3.7) we notice that, from (3.9) and by the definition of $\text{GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$, we know that for a dense set of $\xi\in\mathbb{S}^{n-1}$ conditions (3.7) and (3.8) hold true. From [21, Theorem 3.5], for every such $\xi$ we have
@@
whenever $\tau\in\mathcal{T}$. By exploiting (3.10) and the lower semi-continuity of $\xi\mapsto|D\tau(u\cdot\xi)|(U)$ for every open set $U\subset\Omega$, we argue as before and show that actually (3.10) holds for every $\xi\in\mathbb{S}^{n-1}$ and every Borel set $B\subset\Omega$. By using again [21, Theorem 3.5], we infer that for every $\xi\in\mathbb{S}^{n-1}$ property (3.8) holds true, and also $\hat{u}^{\xi}_{y}\in\text{BV}_{loc}(\Omega^{\xi}_{y})$ for $\mathcal{H}^{n-1}$-a.e. $y\in\Pi^{\xi}$. Eventually, condition (3.8) and the specific form of $\lambda$, together with a simple disintegration argument with respect to the projection $\pi_{\xi}\colon\mathbb{R}^{n}\to\Pi^{\xi}$, give exactly the validity of (3.7). ∎
@@
Definition 3.9 (The outer measure $\mathscr{I}_{u,p}$).
Let $u\colon\Omega\to\mathbb{R}^{n}$ be measurable and let $\{\eta_{\xi}\}_{\xi\in\mathbb{S}^{n-1}}$ be the family of measures in (3.13)–(3.14). Let moreover $f_{B}(\xi):=\eta_{\xi}(B)$ for every $\xi\in\mathbb{S}^{n-1}$ and $B\subset\Omega\text{ Borel}$. For $p\in[1,\infty]$, by virtue of (3.12) we define the set function
$\zeta_{p}(B):=\|f_{B}\|_{L^{p}(\mathbb{S}^{n-1})}.$
Via the classical Caratheodory’s construction we define the (outer) Borel regular measure $\mathscr{I}_{u,p}$ on $\Omega$ as
$\mathscr{I}_{u,p}(E):=\sup_{\delta>0}\,\inf_{G_{\delta}}\sum_{B\in G_{\delta}}\zeta_{p}(B),$

### v2

@@ 2. Definitions and notation
for every measurable function $u\colon B\subset\mathbb{R}\to\mathbb{R}$ such that $A\subset B\cap(B-\varepsilon)$. For the sequel it is convenient to introduce the following class of Mumford-Shah and Griffith functionals. For $\gamma>0$, the former reads as
$\mathcal{MS}_{\gamma}(u,\Omega):=\begin{cases}\gamma\int_{\Omega}|\nabla u(x)|^{2}\textnormal{d}x+\mathcal{H}^{n-1}(J_{u}),\quad u\in\text{GSBV}(\Omega),\\ +\infty,\qquad u\in L^{1}_{loc}(\Omega)\setminus\text{GSBV}(\Omega).\end{cases}$
@@
$\mathcal{G}_{\lambda}(u,\Omega):=\begin{cases}\lambda\int_{\Omega}|e(u)(x)|^{2}\textnormal{d}x+\mathcal{H}^{n-1}(J_{u}),\quad u\in\text{GSBD}(\Omega),\\ \infty,\qquad u\in L^{0}(\Omega;\mathbb{R}^{n})\setminus\text{GSBD}(\Omega).\end{cases}$
@@
Next, we recall a measurability property originally stated in [26, Lemma 3.6].
@@
We say that a $\mathcal{L}^{n}$-measurable function $u\colon\Omega\to\mathbb{R}^{n}$ belongs to the space $\textnormal{GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ if for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ and for $\mathcal{H}^{n-1}$-a.e. $y\in\Pi^{\xi}$ the function $\hat{u}_{y}^{\xi}\in\textnormal{BV}_{loc}(\Omega_{y}^{\xi})$ and
$\int_{\mathbb{S}^{n-1}}\bigg{(}\int_{\Pi^{\xi}}|D\hat{u}_{y}^{\xi}|(\Omega_{y}^{\xi}\setminus J^{1}_{\hat{u}_{y}^{\xi}})+\mathcal{H}^{0}(\Omega_{y}^{\xi}\cap J^{1}_{\hat{u}_{y}^{\xi}})\textnormal{d}\mathcal{H}^{n-1}(y)\bigg{)}\textnormal{d}\mathcal{H}^{n-1}(\xi)<\infty.$ (2.2)
Moreover, we say that $u\in\textnormal{GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ if $u\in\textnormal{GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ and $\hat{u}_{y}^{\xi}\in\textnormal{SBV}_{loc}(\Omega_{y}^{\xi})$ for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ and for $\mathcal{H}^{n-1}$-a.e. $y\in\Pi^{\xi}$.
@@
Let $\Omega\subset\mathbb{R}^{n}$ open and let $u\colon\Omega\to\mathbb{R}^{n}$ be $\mathcal{L}^{n}$-measurable. Then for every $\xi\in\mathbb{S}^{n-1}$ we define the Borel regular measure $\hat{\mu}^{\xi}_{u}$ in $\Omega$ as follows. We define $\hat{\mu}^{\xi}_{u}$ on $\Omega$ as the unique Borel regular measure on $\Omega$ that satisfies for every open set $U\subset\Omega$
$\hat{\mu}^{\xi}_{u}(U):=\int_{\Pi^{\xi}}|D\hat{u}^{\xi}_{y}|(U^{\xi}_{y}\setminus J^{1}_{\hat{u}^{\xi}_{y}})+\mathcal{H}^{0}(U^{\xi}_{y}\cap J^{1}_{\hat{u}^{\xi}_{y}})\,\mathrm{d}\mathcal{H}^{n-1}(y),$
whenever for $\mathcal{H}^{n-1}$-a.e. $y\in\Pi^{\xi}$ the function $\hat{u}^{\xi}_{y}$ belongs to ${\rm BV}_{loc}(U^{\xi}_{y})$, while $\hat{\mu}^{\xi}_{u}(U)=\infty$ otherwise.
@@
Since the hypothesis of De Giorgi-Letta’s Theorem [4] are fulfilled, the extension of the set function $U\mapsto\hat{\mu}^{\xi}_{u}(U)$ to a Borel regular measure on $\Omega$ is uniquely guaranteed by the formula
$\hat{\mu}^{\xi}_{u}(B):=\inf\{\hat{\mu}^{\xi}_{u}(U):B\subset U,\ U\subset\Omega\text{ open}\}.$
@@
In view of (2.2), if $u\in{\rm GBV}^{\mathcal{E}}(\Omega)$ we have that $\hat{\mu}^{\xi}_{u}\in\mathcal{M}^{+}_{b}(\Omega)$ for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$. Moreover, for such $\xi$, we can make use of the measurability property contained in Lemma 2.2 to infer
$\hat{\mu}^{\xi}_{u}(B)=\int_{\Pi^{\xi}}|D\hat{u}^{\xi}_{y}|(B^{\xi}_{y}\setminus J^{1}_{\hat{u}^{\xi}_{y}})+\mathcal{H}^{0}(B^{\xi}_{y}\cap J^{1}_{\hat{u}^{\xi}_{y}})\,\mathrm{d}\mathcal{H}^{n-1}(y),$
@@
For $p\geq 1$, $u\in{\rm GBV}^{\mathcal{E}}(\Omega)$, and $\xi\in\mathbb{S}^{n-1}$ such that $\hat{\mu}^{\xi}_{u}\in\mathcal{M}^{+}_{b}(\Omega)$, we define $\hat{\mu}^{p}_{u}$ on $\Omega$ as the unique Borel regular measure on $\Omega$ that satisfies
$\hat{\mu}^{p}_{u}(U):=\sup_{\mathscr{B}}\sum_{B\in\mathscr{B}}\bigg{(}\int_{\mathbb{S}^{n-1}}\hat{\mu}^{\xi}_{u}(B)^{p}\,\mathrm{d}\mathcal{H}^{n-1}(\xi)\bigg{)}^{\frac{1}{p}}\qquad\text{for $U\subset\Omega$ open,}$, (2.3)
@@
Observe that for $p=1$ the following explicit representation of $\hat{\mu}^{1}_{u}$ holds true
$\hat{\mu}^{1}_{u}(U)=\int_{\mathbb{S}^{n-1}}\hat{\mu}^{\xi}_{u}(U)\,\mathrm{d}\mathcal{H}^{n-1}(\xi)\qquad\text{for $U\subset\Omega$ open}.$ (2.4)
Remark 2.9.
We observe that, by virtue of Jensen’s inequality, it holds true that
$\mu^{1}_{u}\leq\mathcal{H}^{n-1}(\mathbb{S}^{n-1})^{\frac{p-1}{p}}\mu^{p}_{u},$
in the sense of measures for every $p\geq 1$.
@@ 3. Structure of the space $\textnormal{GBV}^{\mathcal{E}}$
The present section addresses the study of the space ${\rm GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ introduced in Definition 2.3. In particular, we focus on the structure of functions $u\in{\rm GSBV}^{\mathcal{E}}(\Omega,\mathbb{R}^{n})$. As a byproduct, we prove the following identification Theorem.
@@
Let $\Omega\subset\mathbb{R}^{n}$ open. A measurable function $u\colon\Omega\to\mathbb{R}^{n}$ belongs to ${\rm GSBD}(\Omega)$ if and only if $u\in{\rm GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$.
@@ 3.1. Preliminaries
We discuss here the measurability of $\xi\mapsto\hat{\mu}^{\xi}_{u}(B)$ for $B$ Borel subset of $\Omega$.
@@
Whenever $\xi\in\mathbb{S}^{n-1}$ and $U\subset\Omega$ are such that $\hat{\mu}^{\xi}_{u}(U)<\infty$, equality (3.3) holds in view of [26, Theorem 3.8]. It remains to consider the case $\hat{\mu}^{\xi}_{u}(U)=\infty$. For simplicity of notation, we denote by $\eta$ the set function given by the right-hand side of (3.3). By contradiction, assume that $\eta(U)<\infty$. In particular, for every $\tau\in\mathcal{T}$ we have
@@
Hence, $D_{\xi}\tau(u\cdot\xi)\in\mathcal{M}_{b}(U)$. By a Carathéodory construction using $\eta$ as Gauge function and the family of all open balls contained in $U$ as set of generators, we obtain a Borel regular measure $\lambda$ on $U$, with $\lambda_{\delta}$ as approximating measure for $\delta>0$ (see, e.g., [29, Section 2.10.1]). For every $V\Subset U$, for every $\delta>0$ small enough, and for every covering $\mathcal{G}$ of $V$ made of open balls in $\mathbb{R}^{n}$ with diameter smaller than $\delta$, we may assume that $\bigcup_{A\in\mathcal{G}}A\subset U$. By Besicovitch covering theorem, there exists a dimensional constant $c(n)>0$ and $\mathcal{G}_{1},\ldots,\mathcal{G}_{c(n)}$ disjoint countable subfamilies of $\mathcal{G}$ such that
@@
In particular, we have that $\lambda_{\delta}(V)\leq c(n)\eta(U)$, which yields the inequality $\lambda(V)\leq c(n)\eta(U)$ for every $V\Subset U$. Taking the limit $V\nearrow U$ we conclude that $\lambda$ is a positive bounded Radon measure on $U$. By using (3.4) it is not difficult to show that actually for every Borel set $B\subset U$ we have that $|D_{\xi}\tau(u\cdot\xi)|(B)\leq\lambda(B)$ whenever $\tau\in\mathcal{T}$. Thus, we are in a position to apply [26, Theorem 3.5] to deduce that $\hat{\mu}^{\xi}_{u}(U)\leq\lambda(U)<\infty$, which is a contradiction. Thus, it must be $\eta(U)=\infty$ and equality (3.3) is satisfied. ∎
@@
Under the assumptions of Proposition 3.3, for every open set $U\subset\Omega$ the map $\xi\mapsto\hat{\mu}^{\xi}_{u}(U)$ is lower-semicontinuous on $\mathbb{S}^{n-1}$.
@@
We notice that equality (3.3) holds for every $U\subset\Omega$ open, for every $\xi\in\mathbb{S}^{n-1}$. In particular, the right-hand side of (3.3) is lower-semicontinuous w.r.t. $\xi\in\mathbb{S}^{n-1}$ for fixed $U\subset\Omega$ open, hence $\xi\mapsto\hat{\mu}^{\xi}_{u}(U)$ is also lower-semicontinuous. ∎
@@
$\hat{\mu}^{1}_{u}\bigg{(}\bigcap_{j=1}^{k}U_{j}\bigg{)}=\int_{\mathbb{S}^{n-1}}\hat{\mu}^{\xi}_{u}\bigg{(}\bigcap_{j=1}^{k}U_{j}\bigg{)}\,\textnormal{d}\mathcal{H}^{n-1}(\xi)\,.$
Passing to the limit as $k\to\infty$, by dominated convergence (recall that $\hat{\mu}^{\xi}_{u}(\Omega)\in L^{1}(\mathbb{S}^{n-1})$) we deduce that
$\hat{\mu}^{1}_{u}(K):=\int_{\mathbb{S}^{n-1}}\hat{\mu}^{\xi}_{u}(K)\,\textnormal{d}\mathcal{H}^{n-1}(\xi)\,.$
@@
In this section we prove that if $u\in\textnormal{GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ then $u\in\textnormal{GSBD}(\Omega)$. Such implication is a consequence of the following two theorems, concerning the structure properties of functions in ${\rm GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$.
@@
Let $u\in{\rm GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ and assume that $\hat{\mu}_{u}^{1}$ is a finite measure. Then for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ we have
@@
$\hat{\mu}^{\xi}_{u}(B)\leq\lambda(B),\quad\text{for every $B\subset\Omega$ Borel}.$ (3.8)
@@
$\hat{\mu}^{\xi}_{u}(B)\leq\lambda(B),\quad\text{for every $B\subset\Omega$ Borel}.$ (3.9)
@@
Since $\lambda$ is a finite Borel measure, the above inequality can be further extended to every Borel set $B\subset\Omega$ by outer regularity. This proves (3.8). In order to verify the validity of (3.7) we notice that, from (3.9) and by the definition of $\text{GSBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$, we know that for a dense set of $\xi\in\mathbb{S}^{n-1}$ conditions (3.7) and (3.8) hold true. From [26, Theorem 3.5], for every such $\xi$ we have
@@
whenever $\tau\in\mathcal{T}$. By exploiting (3.10) and the lower semi-continuity of $\xi\mapsto|D\tau(u\cdot\xi)|(U)$ for every open set $U\subset\Omega$, we argue as before and show that actually (3.10) holds for every $\xi\in\mathbb{S}^{n-1}$ and every Borel set $B\subset\Omega$. By using again [26, Theorem 3.5], we infer that for every $\xi\in\mathbb{S}^{n-1}$ property (3.8) holds true, and also $\hat{u}^{\xi}_{y}\in\text{BV}_{loc}(\Omega^{\xi}_{y})$ for $\mathcal{H}^{n-1}$-a.e. $y\in\Pi^{\xi}$. Eventually, condition (3.8) and the specific form of $\lambda$, together with a simple disintegration argument with respect to the projection $\pi_{\xi}\colon\mathbb{R}^{n}\to\Pi^{\xi}$, give exactly the validity of (3.7). ∎
@@
We note that the definition of $\eta_{\xi}$ in (3.13) does not depend on the representative of $u$ in its Lebesgue class.
Definition 3.9 (The outer measure $\mathscr{I}_{u,1}$).
Let $u\colon\Omega\to\mathbb{R}^{n}$ be measurable and let $\{\eta_{\xi}\}_{\xi\in\mathbb{S}^{n-1}}$ be the family of measures in (3.13)–(3.14). By virtue of (3.12), via the classical Caratheodory’s construction, we define the (outer) Borel regular measure $\mathscr{I}_{u,1}$ on $\Omega$ as
$\mathscr{I}_{u,1}(E):=\sup_{\delta>0}\,\inf_{G_{\delta}}\sum_{B\in G_{\delta}}\int_{\mathcal{S}^{n-1}}\eta_{\xi}(B)\textnormal{d}\mathcal{H}^{n-1}(\xi),$